OpenStudy (anonymous):

Lines v and w are perpendicular. If the slope of line v is 1/6 , what is the slope of line w?

OpenStudy (saifoo.khan):

\[m_1 \times m_2 = -1\]

OpenStudy (anonymous):

thats wrong

OpenStudy (saifoo.khan):

insert the value given and solve.

OpenStudy (anonymous):

When two lines are perpendicular, the slopes are negative reciprocals of each other.

OpenStudy (anonymous):

i dont understand

OpenStudy (saifoo.khan):

what's m1 ? it's 1/6 Insert in the formula, \[m_1 \times m_2 = -1\] \[\frac 16 \times m_2 = -1\]Solve for m2 \[m_2 = -6\]

OpenStudy (anonymous):

 The slope of line v will be the negative reciporacal of the slope of line w. saifoo.khan is showing the correct method for algebraic interpretation :)